OpenStudy (gabylovesyou):

@ganeshie8 i forgot how to do this Solve: -2x - 15 = 6x + 9

OpenStudy (anonymous):

add 2x to both sides

OpenStudy (anonymous):

what do you get ?

OpenStudy (gabylovesyou):

i have no idea what im doing.. 2x + -2x - 15 = 6x + 9 + 2x ?

OpenStudy (anonymous):

ok ill explain ok?

OpenStudy (gabylovesyou):

ok..

OpenStudy (anonymous):

-2x - 15 = 6x + 9 Add 2x to both sides -2x + 2x -15 = 6x - 2x + 9 Add 2x to both sides -15 = 8x + 9 <----- this is your answer. Hope this helps :)

OpenStudy (anonymous):

now continue from -15=8x+9 you have to subtract 9 from both sides : -15-9=8x+9-9 -24=8x now divide both sides by 8 -3=x

OpenStudy (anonymous):

Yep :)

OpenStudy (gabylovesyou):

thank you

OpenStudy (anonymous):

Who?

OpenStudy (gabylovesyou):

both

OpenStudy (anonymous):

No problem. If you need anymore help on anything, just tag me @KngTHund3r